A ball is launched from the top of a building 80 m
Answer #1

a.-

To reach the highest point, the final speed should be zero, so:

v_{f} = v_{0} - gt

0 = v_{0} - gt

v_{0} = gt

t = \frac{v_{0}}{g}

t = \frac{(63m/s)sin50}{9.8m/s^{2}}

ANSWER t = 4.92s

b.-

v{_{f}}^{2} = v{_{0}}^{2} - 2gy

v{_{0}}^{2} = 2gy

y = \frac{v{_{0}}^{2}}{2g}

y = \frac{\left [ (63m/s)sin50 \right ]^{2}}{2(9.8m/s^{2})}

y = 118.83m

And the maximum height above the ground is:

y = 118.83m + 80m

ANSWER y = 198.83m

c.-

The time to complete the parable movement is:

t = 2t_{max}

t = 2(4.92s)

t = 9.84s

And the speed at this time is:

v_{f} = v_{0} + gt

v_{f} = 0 + (9.8m/s)(4.92s)

v_{f} = 48.22m/s

And the time from here to hit the ground is:

v_{f} = v_{0} + gt

t = \frac{v_{0}}{g}

t = \frac{48.22m/a}{9.8m/s^{2}}

t = 4.92s

The complete time is:

t = 4.92s + 9.84s

ANSWER t = 14.76s

d.-

The horizontal component of the movment is constant, so:

x = v_{x}t

x = (63m/s)cos50(9.84s)

ANSWER x = 398.48m

e.-

ANSWER: When the ball reaches its maximum height the speed is zero. So, from here, it starts a free fall movement until it hits the ground. So the acceleration at this time is the gravity.

I hope it helps!!